Donna invested her $33,000 bonus and received a total of $970 in interest after one year. If part of the money returned 4% and the remainder 2.25%, then how much did she invest at each rate?
Answer by checkley77(12844)

.04x+.0225(33,000-x)=970
.04x+742.50-.0225x=970
.0175x=970-742.5
.0175x=227.5
x=227.5/.0175
x=13,000 invested @ 4%.
33,000-13,000=20,000 invested @ 2.25%.
Proof:
.04*13,000+20,000*.0225=970
520+450=970
970=970